Circular No. 7058

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998S IN NGC 3877 P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics; and C. Gerardy and R. Fesen, Dartmouth College, write: “Spectra taken with the MDM 2.4-m telescope on Oct. 27 and the 1.5-m Mt. Hopkins telescope by M. Calkins on Nov. 14 show prominent Hα emission with a flux of 3.6 × 10^{−13} erg cm^{−2} s^{−1}. The Hα emission has a full width at the base of 14 400 km/s and three sharp and well resolved peaks at −4650, −400, and +3700 km/s. The Hα line has a very different structure than seen in the ultraviolet emission (cf. IAUC 7047), but the Hα profile is also seen in He I 587.6- and 1083.0-nm and Paβ lines, and may be indicative of a strong interaction with a circumstellar ring or disk. Photometry obtained with the Mt. Hopkins 1.2-m telescope by A. Szentgyorgyi gives the following magnitudes: Nov. 10.6 UT, J = 16.25, H = 14.85, K = 13.44; Nov. 14.5, U = 18.67, B = 19.09, V = 18.65, R = 17.18, I = 17.59 (uncertainties of 0.05 mag). The V − K color of > 5 magnitudes indicates that a large infrared excess has developed, suggestive of dust formation within the ejecta and consistent with the observed blueshifted metal line profiles. A Tillinghast spectrum taken on Nov. 21.5 shows a 15-percent decline in the flux of the redshifted side of Hα emission relative to the blue side, when compared to the Nov. 14 spectrum, implying continuing dust formation.”

SUPERNOVAE 1997ab, 1998S, 1998et B. E. Schaefer and B. Roscherr, Yale University, report on photometry

from three Type-IIn supernovae obtained with the 3.5-m WIYN telescope on Nov. 21.4 UT: “SN 1997ab has B = 20.02, V = 19.92, and R = 18.74. The event was near maximum on 1996 Apr. 11 at B = 14.7 (Hagen et al. 1997, A.Ap. 324, L29), so our observation is 954 days later with an average decline rate of 0.56 mag per 100 days. SN 1997ab provides a near-unique opportunity to observe a supernova at a reasonable brightness at extremely late times. SN 1998S has U = 19.42, B = 19.06, V = 18.68, R = 17.23, and I = 17.73, with the R magnitude perhaps indicative of a bright Hα line. SN 1998et is just past peak, with U = 18.91, B = 20.25, V = 20.19, R = 19.22, and I = 19.38. Galaxy background uncertainties are of order 0.1 mag.”

COMET C/1998 P1 (WILLIAMS) Visual m_{1} estimates by A. Hale, Cloudcroft, NM: Nov. 15.51 UT, 10.1:

(0.2-m reflector; low altitude); 22.51, 9.4 (Hale, 10×70 binoculars).

1998 November 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green